# Variational Bayesian Bow tie Neural Networks with Shrinkage

### v2

@@ Abstract
Despite the dominant role of deep models in machine learning, limitations persist, including overconfident predictions, susceptibility to adversarial attacks, and underestimation of variability in predictions. The Bayesian paradigm provides a natural framework to overcome such issues and has become the gold standard for uncertainty estimation with deep models, also providing improved accuracy and a framework for tuning critical hyperparameters. However, exact Bayesian inference is challenging, typically involving variational algorithms that impose strong independence and distributional assumptions. Moreover, existing methods are sensitive to the architectural choice of the network. We address these issues by constructing a relaxed version of the standard feed-forward rectified neural network, and employing Polya-Gamma data augmentation tricks to render a conditionally linear and Gaussian model. Additionally, we use sparsity-promoting priors on the weights of the neural network for data-driven architectural design. To approximate the posterior, we derive a variational inference algorithm that avoids distributional assumptions and independence across layers and is a faster alternative to the usual Markov Chain Monte Carlo schemes.
@@ 1 Introduction
Neural networks (NNs) are effective deep models that play a dominant role in machine learning and have achieved remarkable success across various domains including medicine and biological sciences Jumper et al. (2021); Yu et al. (2021), natural language processing Mikolov et al. (2013); Touvron et al. (2023), computer vision and image analysis Dosovitskiy et al. (2020), data privacy and security Yang et al. (2019) and beyond. However, modern machine learning applications often lack reliable, if not any, uncertainty estimates Guo et al. (2017); Gal (2016); Ashukha et al. (2020). Classical deep models are easily fooled and are susceptible to adversarial attacks Szegedy et al. (2014); Nguyen et al. (2015); Zong et al. (2024), and even when the adversarial attacks fail, the saliency interpretations of deep neural networks (DNNs) are rather brittle Carbone et al. (2022). When data variations leading to out-of-distribution (OOD) shifts occur neural networks often fail to generalize well Hein et al. (2019); Zhang et al. (2024); Ashukha et al. (2020). Moreover, standard neural networks usually lack intuitive interpretation and explainability and so are regarded as black boxes Lipton (2018). To address these challenges, Bayesian neural networks (BNNs) have emerged as a compelling extension of conventional neural networks (for a review, see e.g. Jospin et al. (2022); Arbel et al. (2023)). While finite (non-Bayesian) deep ensembles of independent neural networks have been shown to improve prediction and uncertainty estimates Lakshminarayanan et al. (2017), the Bayesian approach creates infinite ensembles of deep neural networks. The advantage of this approach is that it controls the model complexity and builds regularization into the model by marginalizing out the parameters. Indeed, Bayesian neural networks have become the gold standard for uncertainty estimation in the context of data-driven decision-making and in safety-critical applications, where robustness and calibration are crucial McAllister et al. (2017); Carbone et al. (2020); Gruver et al. (2023); Yang et al. (2024); Klarner et al. (2023).
@@
A core problem of Bayesian machine learning lies in performing the inference; in practice, the posterior distribution of the model’s parameters given observations is not available in the closed form and direct sampling from the posterior is computationally expensive, meaning one has to employ approximate Bayesian inference. Markov chain Monte Carlo (MCMC) is a gold standard solution since it produces draws, which are asymptotically exact samples from the posterior but for large data sets or complex models with multimodal posteriors it can be prohibitively slow. Variational inference Hinton and van Camp (1993); Jordan et al. (1999) (VI) instead utilizes optimization rather than sampling making it a more computationally effective method suitable for high-dimensional problems. VI approximates the posterior with the closest (most commonly, in terms of the Kullback–Leibler divergence) member of some tractable variational family of distributions taken as close as possible to the true posterior Blei et al. (2017). Recently, several variational algorithms and methods have been proposed and proven to achieve desirable consistency and predictive performance Bai et al. (2020); Zhang et al. (2018); Castillo and Egels (2024); Yang et al. (2020); Chérief-Abdellatif (2020).
@@
In this paper, we follow the setup of Smith et al. (2021) and introduce a bow tie neural network, where a stochastic relaxation of the rectified linear unit (ReLU) activation function leads to a model amendable to the Polya-Gamma (PG) data augmentation trick Polson et al. (2013) and results in conditionally linear and Gaussian stochastic activations. Additionally, on the weights of the network, we place sparsity-inducing priors, which are known for their ability to provide improvement in the predictive performance of Bayesian deep models; not only do sparse models ease the storage and computational burden, they also improve the calibration and may recover the potential sparse structure of the target function Polson and Ročková (2018); Bai et al. (2020); Griffin and Brown (2021); Law and Zankin (2022); Ray and Szabó (2022). Specifically, we consider sparsity-inducing global-local normal-generalized inverse Gaussians (N-GIG) priors Polson and Scott (2012). Section 2 describes the bow tie model with shrinkage priors and implementation of PG data augmentation. Having constructed the bow tie neural network, we propose a (block) structured mean-field family for the approximate variational posterior which is flexible enough and doesn’t require assumptions on the distributional form of each component as well as on independence across layers. For the chosen family, coordinate ascent variation inference (CAVI) Bishop (2016) can be performed, with all variational updates available in the closed form. Whilst continuous shrinkage priors result in more tractable computations, they do not incur exact zeros on the neural network’s weights. We address this issue by implementing a simple post-process node selection algorithm controlled by the empirical Bayesian false discovery rate (FDR). Finally, we derive the predictive distribution and propose improving the accuracy and uncertainty estimation by considering ensembles of variational approximations obtained by running several parallel variational algorithms with different random starting points. In this way, our approach accounts for the multimodality of the posterior distributions arising in Bayesian deep models. Section 3 derives the inference algorithm, variable selection procedure and predictive distribution. We evaluate our method on a range of classical regression tasks as well as synthetic regression tasks and demonstrate its competitiveness compared to alternative well-known Bayesian algorithms in Section 4  [^1] We provide a Python implementation of our model at https://github.com/sheinkmana/V_bowtie_NN..
@@ 2.1 Bow tie neural networks
We begin by describing the class of recently proposed bow tie networks Smith et al. (2021), which are deep generative models that generalize feed-forward rectified linear neural networks with stochastic activations. Let $\mathbf{x}_{n}\in\mathbb{R}^{D_{0}}$ be the inputs, $\mathbf{y}_{n}\in\mathbb{R}^{D_{L+1}}$ be the outputs and $\mathbf{a}_{n}=\{\mathbf{a}_{n,l}\}_{l=1}^{L}$ with $\mathbf{a}_{n,l}\in\mathbb{R}^{D_{l}}$ be the latent activations at each of the $L$ intermediate layers. For notational purposes, assume $\mathbf{a}_{n,0}=\mathbf{x}_{n}$. The model assumes:
@@
Note that Equation 1 is a stochastic relaxation of the standard feed-forward NN, which is recovered in the limiting case when $\bm{\Sigma}_{l}\rightarrow\mathbf{0}$ for $l=1,\ldots,L$. Instead of relying on local gradient-based algorithms, Smith et al. (2021) introduces another relaxation of the model and employs a Polya-Gamma data augmentation trick Polson et al. (2013) to render the model conditionally linear with Gaussian activations. Specifically, consider the ReLU activation function $f(z)=\max(0,z)$. It can alternatively be written as a product of $z$ and a binary function $\gamma$, i.e. $f(z)=\gamma z$ where $\gamma=\textbf{1}(z>0)$. In this way, $\gamma$ determines whether the node is activated ($\gamma=1$) or not ($\gamma=0$). In a similar fashion, the additional stochastic relaxation replaces $f(\mathbf{z}_{n,l})$ with $\bm{\gamma}_{n,l}\odot\mathbf{z}_{n,l}$:
@@
$\gamma_{n,l,d}$ $\overset{ind}{\sim}\text{Bern}\left(\sigma\left(z_{n,l,d}/\mathrm{T}\right)\right),$
where $\mathrm{T}\geq 0$ is the temperature parameter, $\sigma(x)=\exp(x)/(1+\exp(x))$ is the logistic function and $\odot$ represents the elementwise product. Thus, the nodes are turned off or on stochastically depending on their input. Note that after marginalizing over the binary activations, the latent variables $\mathbf{a}_{n}$ are distributed as a mixture of two normals:
$a_{n,l,d}|z_{n,l,d},\theta$ $\sim\sigma(z_{n,l,d}/\mathrm{T})\text{N}\left(z_{n,l,d},\eta_{l,d}^{2}\right)+\left(1-\sigma(z_{n,l,d}/\mathrm{T})\right)\text{N}\left(0,\eta_{l,d}^{2}\right),$ (2)
@@
$\mathbb{E}[a_{n,l,d}|z_{n,l,d},\theta]$ $=\mathbb{E}[\mathbb{E}[a_{n,l,d}|z_{n,l,d},\gamma_{n,l,d},\theta]]=\mathbb{E}[\gamma_{n,l,d}z_{n,l,d}]$
@@
$\mathbb{V}(a_{n,l,d}|z_{n,l,d},\theta)$ $=\mathbb{E}[\mathbb{V}(a_{n,l,d}|z_{n,l,d},\gamma_{n,l,d},\theta)]+\mathbb{V}(\mathbb{E}[a_{n,l,d}|z_{n,l,d},\gamma_{n,l,d},\theta])$
@@
We display the conditional distribution of $a_{n,l,d}$ in Figure 1, for different combinations of the temperature parameter $T$ and variance $\eta_{l,d}^{2}$. The ReLU activation is recoverd in the case of $T=0$ and $\eta_{l,d}^{2}=0$, while other choices of $T$ and $\eta_{l,d}^{2}$ generalize the ReLU. The density resembles a bow tie, hence the name of the model.
@@ 2.2 Shrinkage Priors
Prior elicitation in Bayesian neural networks is challenging, as understanding how the high-dimensional weights map to the functions implemented by the network is not trivial. Standard Gaussian priors are often a default choice, also due to their link with $\ell_{2}$ regularization in maximum a posteriori (MAP) inference; indeed, such priors were used in Smith et al. (2021). For an overview and discusson on priors in Bayesian neural networks, see Fortuin (2022).
We take an alternative approach to the Gaussian priors of Smith et al. (2021) in order to sparsify our model. Sparsity-inducing priors ease the problem of storage and computational costs, have been shown to provide improvement in the predictive performance of deep models, and can provide a data-driven approach to selecting the width and depth, easing the difficult task of specifying the network architecture. For these reasons, such priors have been considered in a large number of works to produce sparse BNNs. For classical proposals on the two-group discrete mixture priors with a point mass at zero (referred to as spike-and-slab priors) in high-dimensional regression, see George and McCulloch (1993); Mitchell and Beauchamp (1988). More recently, many consider spike-and-slab priors on the neural network weights and provide both theoretical guarantees and demonstrate empirical improvements Blundell et al. (2015); Polson and Ročková (2018); Bai et al. (2020); Sun et al. (2022); Lee and Lee (2022). As an alternative to spike-and-slab priors, shrinkage priors employ a single distribution to approximate the spike-and-slab shape, yet are more computationally attractive, as they avoid exploring the space of all possible models (corresponding to selecting subsets of inputs/nodes). Moreover, (nearly) optimal theoretical guarantees of spike-and-slab priors can still be obtained with shrinkage priors in the linear regression setting Song and Liang (2023). A popular choice of shrinkage prior is the horseshoe Carvalho et al. (2009); Piironen and Vehtari (2017), and imposing regularized horseshoe priors on the BNN weights combined with a structured variational approximation provides competitive empirical results Ghosh et al. (2018). Another example arising in neural networks is a Gaussian scale mixture prior with automatic relevance determination, which when combined with suitable approximate Bayesian inference is equivalent to introducing dropout regularization in NN Nalisnick et al. (2019). Recently, Castillo and Egels (2024) shows that a suitably rescaled heavy-tailed prior on the neural network weights achieves automatic adaptation, simultaneously to both the intrinsic dimension and smoothness of the underlying function, and near-optimal minimax contraction rates of the fractional posterior distribution and its mean-field variational approximation.
In this work, we focus on a class of continuous shrinkage priors, namely, global-local normal scale-mixtures with generalized inverse Gaussian shrinkage priors on the scale parameters, referred to as global-local normal-generalized inverse Gaussian priors Griffin and Brown (2021). Global-local scale-mixtures aim to shrink less important weights whilst leaving large ones, which is achieved through a global parameter controlling the overall shrinkage, with the local parameters allowing deviations at the level of individual nodes Polson and Scott (2012); Bhadra et al. (2019). This choice of priors is also motivated by the theoretical guarantees for high-dimensional regression Song and Liang (2023); Griffin and Brown (2010); Polson et al. (2013), for a survey on global-local shrinkage methods we refer to Griffin and Brown (2021).
@@
with parameters $\nu$, $\delta$, and $\lambda$; for a proper prior, $\nu>0$ if $\delta=0$ or $\nu<0$ if $\lambda=0$. In Equation 6, we allow the GIG parameters for the local scale parameters $\psi_{l,d,d^{\prime}}$ to vary across layers to adjust local shrinkage for wider layers. Furthermore, to encourage more shrinkage for larger depth and width, we scale the global parameters $\tau_{l}$ with respect to $L$ and the local parameters $\psi_{l,d,d^{\prime}}$ with respect to $D_{l}$ (details of our approach are provided Section C.1).
When the global shrinkage parameter $\tau_{l}$ is fixed, examples of the marginal distribution for $w_{l,d,d^{\prime}}$ include Laplace Park and Casella (2008), Student-t (ST) Tipping (2001), Normal-Gamma (NG) Caron and Doucet (2008); Griffin and Brown (2010), Normal inverse Gaussian (NIG) Caron and Doucet (2008). Each example has a different tail behavior, inducing different forms of shrinkage (see Table 1 for a overview and Figure 2 (a) for a visualization). Note that if the prior is polynomial-tailed, then for large signals the amount of shrinkage is mitigated even given small $\tau_{l}$ Polson and Scott (2012).
@@
Mixing distribution IG Gamma Gamma IGauss
@@
The global shrinkage parameter $\tau_{l}$ leads to a non-separable penalty for the weights within the same layer, i.e. after integrating out $\tau_{l}$, the weights within the same layer are dependent. This is illustrated in Figure 2 (b), which provides contour plots for the joint marginal distribution of two weights, within the same layer (dependence, left column) and across different layers (independence, right column), for the two choice of IG and Gamma mixing priors. Figure 2 (b) also highlights how the variance depends on the width of the layer, with more hidden units and smaller variance for the second layer compared to the first.
@@
Figure 2: Illustration of the prior on the weights. (a) the marginal density of the weights for different choices within the GIG family. (b) the joint prior on two weights within the same layer (left) and across layers (right) for the two choices of IG (top) and Gamma (bottom) mixing priors.
The opposite effects of varying width and depth in deep neural networks are studied in Vladimirova et al. (2021); while depth accentuates a model’s non-Gaussianity, the width makes models increasingly Gaussian. Indeed, infinitely wide BNNs are closely related to Gaussian processes (GPs), typically relying on appropriately scaled i.i.d. Gaussian weights Neal (2012); Lee et al. (2018); Matthews et al. (2018) and relaxing these assumptions, e.g. through ordering, constraints, heavy tails, or bottlenecks, results in non-Gaussian limits, such as stable processes Peluchetti et al. (2020), deep GPs Agrawal et al. (2020) or more exotic processes Sell and Singh (2023); Chada et al. (2022). The sparsity promoting priors in Equations 5, 6 and 7 provide a framework for the data to inform on the width and depth of the network.

### v3

@@ Abstract
Despite the dominant role of deep models in machine learning, limitations persist, including overconfident predictions, susceptibility to adversarial attacks, and underestimation of variability in predictions. The Bayesian paradigm provides a natural framework to overcome such issues and has become the gold standard for uncertainty estimation with deep models, also providing improved accuracy and a framework for tuning critical hyperparameters. However, exact Bayesian inference is challenging, typically involving variational algorithms that impose strong independence and distributional assumptions. Moreover, existing methods are sensitive to the architectural choice of the network. We address these issues by focusing on a stochastic relaxation of the standard feed-forward rectified neural network and using sparsity-promoting priors on the weights of the neural network for increased robustness to architectural design. Thanks to Polya-Gamma data augmentation tricks, which render a conditionally linear and Gaussian model, we derive a fast, approximate variational inference algorithm that avoids distributional assumptions and independence across layers. Suitable strategies to further improve scalability and account for multimodality are considered.
@@ 1 Introduction
Neural networks (NNs) are effective deep models that play a dominant role in machine learning and have achieved remarkable success across various domains including medicine and biological sciences (Jumper et al., 2021; Yu et al., 2021), natural language processing (Mikolov et al., 2013; Touvron et al., 2023), computer vision and image analysis (Dosovitskiy et al., 2020), data privacy and security (Yang et al., 2019) and beyond. However, modern machine learning applications often lack reliable, if any, uncertainty estimates (Guo et al., 2017; Gal, 2016; Ashukha et al., 2020). Classical deep models are easily fooled and are susceptible to adversarial attacks (Szegedy et al., 2014; Nguyen et al., 2015; Zong et al., 2024), and even when the adversarial attacks fail, the saliency interpretations of deep neural networks (DNNs) are rather brittle (Carbone et al., 2022). When data variations leading to out-of-distribution (OOD) shifts occur, neural networks often fail to generalize well (Hein et al., 2019; Zhang et al., 2024; Ashukha et al., 2020). Moreover, standard neural networks usually lack intuitive interpretation and explainability and so are regarded as black boxes (Lipton, 2018). To address these challenges, Bayesian neural networks (BNNs) have emerged as a compelling extension of conventional neural networks (for a review, see e.g. (Jospin et al., 2022; Arbel et al., 2023)). While finite (non-Bayesian) deep ensembles of independent neural networks have been shown to improve prediction and uncertainty estimates (Lakshminarayanan et al., 2017), the Bayesian approach creates infinite ensembles of deep neural networks. The advantage of this approach is that it controls the model complexity and builds regularization into the model by marginalizing the parameters. Indeed, BNNs have become the gold standard for uncertainty estimation in the context of data-driven decision-making and in safety-critical applications, where robustness and calibration are crucial (McAllister et al., 2017; Carbone et al., 2020; Gruver et al., 2023; Yang et al., 2024; Klarner et al., 2023).
@@
A core problem of Bayesian machine learning lies in performing inference; in practice, the posterior distribution of the model’s parameters given observations is not available in closed form, and direct sampling from the posterior is computationally expensive, meaning one has to employ approximate Bayesian inference. Markov chain Monte Carlo (MCMC) is a gold standard solution since it produces draws, which are asymptotically exact samples from the posterior, but for large data sets or complex models with multimodal posteriors, it can be prohibitively slow. Variational inference (VI Jordan et al., 1999; Blei et al., 2017) instead utilizes optimization rather than sampling making it a more computationally effective method suitable for high-dimensional, large-scale problems. VI approximates the posterior with the closest (most commonly, in terms of the Kullback–Leibler divergence) member of some tractable variational family of distributions taken as close as possible to the true posterior.
In this work, we focus on advancing approximate inference for BNNs. Specifically, we improve the bow tie neural network of (Smith et al., 2021) by introducing sparsity-inducing priors, constructing a fast, approximate variational inference algorithm, and exploring further strategies to improve computational gains and performance in terms of both accuracy and uncertainty estimation. Below, we review related research and provide an outline of the paper and its contributions.
Related work.
Sparsity-inducing priors are known for their ability to improve robustness to overparametrization in Bayesian modeling; they also lead to better calibrated uncertainty and, in certain settings, may recover the sparse structure of the target function (e.g. Castillo et al., 2015; Song and Liang, 2023; Song, 2020). Such priors generally fall within two classes: 1) the two-group discrete mixture priors with a point mass at zero (referred to as spike-and-slab priors) (George and McCulloch, 1993; Mitchell and Beauchamp, 1988) or 2) shrinkage priors, which employ a single distribution to approximate the spike-and-slab shape, yet are more computationally attractive, as they avoid exploring the space of all possible models. Both types of priors have become widely used in Bayesian deep modeling, due to their high-dimensionality and overparametrization, and are further supported by theoretical guarantees ((Polson and Ročková, 2018; Sun et al., 2022) for BNNs with spike-and-slab priors and (Castillo and Egels, 2024; Lee and Lee, 2022) for BNNs with heavy-tailed shrinkage priors).
To scale with the size of the data and model complexity, various variational algorithms and methods have been proposed for sparse BNNs (for an overview, see Zhang et al., 2018). Specifically, variational inference techniques for BNNs with spike-and-slab priors include (Blundell et al., 2015; Bai et al., 2020b, a; Jantre et al., 2024). Within the class of shrinkage priors, horseshoe priors (Carvalho et al., 2009; Piironen and Vehtari, 2017a) on the BNN weights combined with variational approximations have been shown to provide competitive empirical results (Louizos et al., 2017; Ghosh et al., 2018, 2019). Dropout regularization in neural networks has also been shown to be closely connected to sparse BNNs with suitable variational approximations (Kingma et al., 2015; Hron et al., 2018; Louizos et al., 2017; Liu et al., 2019; Nalisnick et al., 2019). In a slightly different direction, (Li et al., 2024) consider BNNs with Gaussian priors and introduce a variational inference methodology which enforces sparsity in the weights during the training. Finally, we mention the adaptive variational Bayes method of (Ohn and Lin, 2024), which was successfully applied to Bayesian deep learning scenarios; the framework operates on a collection of models, considers "sieve" priors (Arbel et al., 2013) to combine several variational approximations.
Contributions and outline.
@@
(Smith et al., 2021) introduced a bow tie neural network, where a stochastic relaxation of the rectified linear unit (ReLU) activation function leads to a model amenable to the Polya-Gamma (PG) data augmentation trick (Polson et al., 2013) and results in conditionally linear and Gaussian stochastic activations. In this paper, we improve bow tie neural networks in several ways. First, to improve robustness with increasing width and depth of the networks, we place sparsity-inducing global-local normal-generalized inverse Gaussians (N-GIG) priors (Polson and Scott, 2012) on the weights of the network. Section 2 describes the bow tie model with shrinkage priors and implementation of PG data augmentation. Second, while (Smith et al., 2021) focus on MCMC, in Section 3 we propose a (block) structured mean-field family for the approximate variational posterior, which is flexible and doesn’t require parametric assumptions on the distributional form of each component as well as on independence across layers. For the chosen family, fast coordinate ascent variation inference (CAVI) (Bishop, 2016) can be performed, with all variational updates available in the closed form. Third, to improve the scalability of the algorithm, we consider two strategies: a stochastic variant that employs subsampling to cope with large data and a post-process node selection algorithm to obtain a sparse posterior that eases the storage and computational burden of predictions. Fourth, we propose improving accuracy and uncertainty estimation by considering ensembles of variational approximations obtained by running several parallel variational algorithms with different random starting points. In this way, our approach accounts for the multimodality of the posterior distributions arising in Bayesian deep models. Sections 3.1 and 3.2 derive the inference algorithm and in Section 3.3 a stochastic variant (Hoffman et al., 2013) of the algorithm is developed. Further, we derive a variable selection procedure in Section 3.4 for faster prediction (Section 3.5) and propose ensembles in Section 3.6. We evaluate our method [^1] We provide a Python implementation of our model on GitHub. on a range of classical regression tasks as well as synthetic regression tasks and demonstrate its competitiveness compared to alternative, well-known Bayesian algorithms in Section 4.
@@ 2.1 Bow tie Neural Networks
We begin by describing the class of recently proposed bow tie networks (Smith et al., 2021), which are deep generative models that generalize feed-forward rectified linear neural networks with stochastic activations. Let $\mathbf{x}_{n}\in\mathbb{R}^{D_{0}}$ be the inputs, $\mathbf{y}_{n}\in\mathbb{R}^{D_{L+1}}$ be the outputs and $\mathbf{a}_{n}=\{\mathbf{a}_{n,l}\}_{l=1}^{L}$ with $\mathbf{a}_{n,l}\in\mathbb{R}^{D_{l}}$ be the latent activations at each of the $L$ intermediate layers. For notational purposes, assume $\mathbf{a}_{n,0}=\mathbf{x}_{n}$. The model assumes:
@@
Note that Equation 1 is a stochastic relaxation of the standard feed-forward NN, which is recovered in the limiting case when $\bm{\Sigma}_{l}\rightarrow\mathbf{0}$ for $l=1,\ldots,L$. Instead of relying on local gradient-based algorithms, (Smith et al., 2021) introduces another relaxation of the model and employs a Polya-Gamma data augmentation trick (Polson et al., 2013) to render the model conditionally linear with Gaussian activations. Specifically, consider the ReLU activation function $f(z)=\max(0,z)$. It can alternatively be written as a product of $z$ and a binary function $\gamma$, i.e. $f(z)=\gamma z$ where $\gamma=\textbf{1}(z>0)$. In this way, $\gamma$ determines whether the node is activated ($\gamma=1$) or not ($\gamma=0$). The additional stochastic relaxation assumes that $\gamma$ is not deterministic, but a binary random variable whose success probability depends on $z$. Specifically, the dependence is constructed through the logistic function $\sigma(x)=\exp(x)/(1+\exp(x))$ and a temperature parameter $\mathrm{T}\geq 0$:
@@
$\gamma_{n,l,d}$ $\overset{ind}{\sim}Bern\left(\sigma\left(z_{n,l,d}/\mathrm{T}\right)\right)\text{ for }T>0,$
where $\odot$ represents the elementwise product. Thus, the nodes are turned off or on stochastically depending on their input. Note that in the limit as the temperature $T\rightarrow 0$, we have that $\gamma_{n,l,d}=\textbf{1}(z_{n,l,d}>0)$ and $\mathbf{a}_{n,l}|\mathbf{z}_{n,l},\bm{\gamma}_{n,l},\bm{\theta}\sim\text{N}\left(\textbf{1}(\mathbf{z}_{n,l}>0)\odot\mathbf{z}_{n,l},\bm{\Sigma}_{l}\right)$.
For $T>0$, after marginalizing the binary activations, the stochastic activations $\mathbf{a}_{n}$ are distributed as a mixture of two normals:
${\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}|z_{n,l,d},\bm{\theta}$ $\sim\sigma(z_{n,l,d}/\mathrm{T})\text{N}\left(z_{n,l,d},\eta_{l,d}^{2}\right)+\left(1-\sigma(z_{n,l,d}/\mathrm{T})\right)\text{N}\left(0,\eta_{l,d}^{2}\right),$ (2)
@@
$\mathbb{E}[{\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}|z_{n,l,d},\bm{\theta}]$ $=\mathbb{E}[\mathbb{E}[{\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}|z_{n,l,d},\gamma_{n,l,d},\bm{\theta}]]=\mathbb{E}[\gamma_{n,l,d}z_{n,l,d}]$
@@
$\mathbb{V}({\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}|z_{n,l,d},\bm{\theta})$ $=\mathbb{E}[\mathbb{V}({\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}|z_{n,l,d},\gamma_{n,l,d},\bm{\theta})]+\mathbb{V}(\mathbb{E}[{\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}|z_{n,l,d},\gamma_{n,l,d},\bm{\theta}])$
@@
The conditional distribution of ${\color[rgb]{0,0,0}{\mathrm{a}_{n,l,d}}}$ is displayed in Figure 1, for different combinations of the temperature parameter $T$ and variance $\eta_{l,d}^{2}$. The ReLU activation is recovered in the case of $T=0$ and $\eta_{l,d}^{2}=0$, while other choices of $T$ and $\eta_{l,d}^{2}$ generalize the ReLU. The stochastic activations have smoother conditional mean functions for non-zero values of $T$, with potential bimodality (density resembles a bow tie, hence the name) for larger $T$ relative to $\eta$. This may be relevant in settings with multimodal predictive distributions, although our experiments focus on small temperature values for more similarity to the ReLU.
@@ 2.2 Shrinkage Priors
Prior elicitation in Bayesian neural networks is challenging, as understanding how the high-dimensional weights map to the functions implemented by the network is not trivial. Standard Gaussian priors are often a default choice, also due to their link with $\ell_{2}$ regularization in maximum a posteriori (MAP) inference; indeed, such priors were used in (Smith et al., 2021). For an overview and discussion on priors in Bayesian neural networks, see (Fortuin, 2022).
We take an alternative approach to the Gaussian priors of (Smith et al., 2021) in order to sparsify our model. Sparsity-inducing priors have been shown to provide improvement in the predictive performance of deep models and can provide a data-driven approach to selecting the width and depth, easing the difficult task of specifying the network architecture. In this work, we focus on a class of continuous shrinkage priors, namely, global-local normal scale-mixtures with generalized inverse Gaussian shrinkage priors on the scale parameters, referred to as global-local normal-generalized inverse Gaussian priors (Griffin and Brown, 2021). Global-local scale-mixtures aim to shrink less important weights whilst leaving large ones, which is achieved through a global parameter controlling the overall shrinkage, with the local parameters allowing deviations at the level of individual nodes (Polson and Scott, 2012; Bhadra et al., 2019). This choice of priors is also motivated by the theoretical guarantees for high-dimensional regression (Song and Liang, 2023; Griffin and Brown, 2010; Polson et al., 2013), for a survey on global-local shrinkage methods we refer to (Griffin and Brown, 2021).
@@
with parameters $\nu$, $\delta$, and $\lambda$; for a proper prior, $\nu>0$ if $\delta=0$ or $\nu<0$ if $\lambda=0$. In Equation 6, we allow the GIG parameters for the local scale parameters $\psi_{l,d,d^{\prime}}$ to vary across layers to adjust local shrinkage for wider layers. Furthermore, to encourage more shrinkage for larger depth and width, we scale the global parameters $\tau_{l}$ with respect to $L$ and the local parameters $\psi_{l,d,d^{\prime}}$ with respect to $D_{l}$ (details of our approach are provided Section D.1).
When the global shrinkage parameter $\tau_{l}$ is fixed, examples of the marginal distribution for $w_{l,d,d^{\prime}}$ include Laplace (Park and Casella, 2008), Student-t (ST) (Tipping, 2001), Normal-Gamma (NG) (Caron and Doucet, 2008; Griffin and Brown, 2010), Normal inverse Gaussian (NIG) (Caron and Doucet, 2008). Each example has a different tail behavior, inducing different forms of shrinkage (see Table 1 for an overview and Figure 2 (a) for a visualization). Note that if the prior is polynomial-tailed, then for large signals the amount of shrinkage is mitigated even given small $\tau_{l}$ (Polson and Scott, 2012).
@@
Mixing distribution Inverse Gamma (IG) Gamma Gamma Inverse Gaussian (IGauss)
@@
The global shrinkage parameter $\tau_{l}$ leads to a non-separable penalty for the weights within the same layer, i.e. after integrating out $\tau_{l}$, the weights within the same layer are dependent. This is illustrated in the top row of Figure 2 (b), which shows how the conditional prior density of one weight varies given different values of another weight within the same layer, for two choices of inverse Gamma (IG) and Gamma mixing priors. Instead, across layers the weights are independent, as illustrated by the contour plots for the joint density in the bottom row of Figure 2 (b). The bottom row of Figure 2 (b) also highlights how the variance depends on the width of the layer, with more hidden units and smaller variance for the second layer compared to the first.
@@
Figure 2: Illustration of the prior on the weights. (a) the marginal density of the weights for different choices within the GIG family. (b) the conditional prior of the weights within the same layer (top) and joint prior of the weights across layers (bottom) for two choices of IG (left) and Gamma (right) mixing priors.
The opposite effects of varying width and depth in deep neural networks are studied in (Vladimirova et al., 2021); while depth accentuates a model’s non-Gaussianity, the width makes models increasingly Gaussian. Indeed, infinitely wide BNNs are closely related to Gaussian processes (GPs), typically relying on appropriately scaled i.i.d. Gaussian weights (Neal, 2012; Lee et al., 2018; Matthews et al., 2018) and relaxing these assumptions, e.g. through ordering, constraints, heavy tails, or bottlenecks, results in non-Gaussian limits, such as stable processes (Peluchetti et al., 2020), deep GPs (Agrawal et al., 2020) or more exotic processes (Sell and Singh, 2023; Chada et al., 2022). The sparsity-promoting priors in Equations 5, 6 and 7 provide a framework for the data to inform on the width and depth of the network.